### v2

# Trace field degrees in the Torelli group

Abstract.
We show that for $g\geq 2$, all integers $1\leq d\leq 3g-3$ arise as trace field degrees of pseudo-Anosov mapping classes in the Torelli group of the closed orientable surface of genus $g$. Our method uses the Thurston–Veech construction of pseudo-Anosov maps, and we provide examples where the stretch factor has algebraic degree any even number between two and 6g66g-6. This validates a claim by Thurston from the 1980s.
1. Introduction
Let $S_{g}$ be the closed orientable surface of genus $g\geq 2$. A homeomorphism $f$ of $S_{g}$ is pseudo-Anosov if there exists a pair of transverse singular measured foliations $\mathcal{F}^{u}$ and $\mathcal{F}^{s}$ and a real number $\lambda>1$ such that $f(\mathcal{F}^{u})=\lambda\mathcal{F}^{u}$ and $f(\mathcal{F}^{s})=\lambda^{-1}\mathcal{F}^{s}$. The number $\lambda>1$ is the stretch factor of the pseudo-Anosov map, and is an algebraic integer. The degree of the field extension $\mathbb{Q}(\lambda):\mathbb{Q}$ is the stretch factor degree, and is bounded from above by the dimension of the Teichmüller space of $S_{g}$, namely 6g66g-6 [Thu88].
There is another field which plays a central role in this article: $\mathbb{Q}(\lambda+\lambda^{-1})$. This field is the trace field and is uniquely determined by the pair $(\mathcal{F}^{u},\mathcal{F}^{s})$ [KS00, GJ00]. The degree of the field extension $\mathbb{Q}(\lambda+\lambda^{-1}):\mathbb{Q}$ is the trace field degree, and is bounded from above by 3g33g-3.
Strenner showed that for $S_{g}$, all integers $1\leq d\leq 3g-3$ arise as trace field degrees of pseudo-Anosov maps [Str17]. Furthermore, Strenner determined the set of integers which arise as stretch factor degrees: all even integers between $2$ and $\leq 6g-6$, as well as all odd integers between $3$ and 3g33g-3 [Str17], but none of Strenner’s examples are elements of the Torelli group (see Section 1.3).
1.1. Torelli groups
The Torelli group $\mathcal{I}(S_{g})$ is the kernel of the symplectic representation of the mapping class group $\mathrm{Mod}(S_{g})$. In [Mar19, Problem 10.6], Margalit asked which stretch factor degrees arise for pseudo-Anosov mapping classes in the Torelli group. Our first main result completely answers the question of trace field degrees arising in Torelli groups.
Theorem 1.
Let $g\geq 2$. Every integer $1\leq d\leq 3g-3$ arises as the trace field degrees of a pseudo-Anosov mapping class in the Torelli group $\mathcal{I}(S_{g})$.
The stretch factor $\lambda$ satisfies the quadratic equation $t^{2}-(\lambda+\lambda^{-1})t+1$ over the trace field $\mathbb{Q}(\lambda+\lambda^{-1})$. Hence, the field extension $\mathbb{Q}(\lambda):\mathbb{Q}(\lambda+\lambda^{-1})$ has degree one or two. We obtain our second main result by showing that for all trace field degrees there exist instances where this field extension degree equals two.
Theorem 2.
Let $g\geq 2$. Every even integer $2d$ with $2\leq 2d\leq 6g-6$ arises as the stretch factor degree of a pseudo-Anosov mapping class in the Torelli group $\mathcal{I}(S_{g})$.
1.2. The Thurston–Veech construction
We prove our main results using the Thurston–Veech construction. This construction of pseudo-Anosov maps appeared independently in two papers by Thurston and Veech [Thu88, Vee89].
A multicurve is a disjoint union of simple closed curves, and a pair of multicurves $\alpha,\beta\subset S_{g}$ fills the surface $S_{g}$ if $\alpha$ and $\beta$ intersect transversally and if the complement $S_{g}\setminus\left(\alpha\cup\beta\right)$ is a union of topological discs none of which is a bigon. This in particular implies that each pair $\alpha_{i}$ and $\beta_{j}$ of components realises the minimal number of intersection points within their respective isotopy classes.
Given a pair of filling multicurves $\alpha,\beta\subset S_{g}$, the Thurston–Veech construction provides pseudo-Anosov mapping classes in the subgroup $\langle T_{\alpha},T_{\beta}\rangle$ of $\mathrm{Mod}(S_{g})$ generated by multitwists along the multicurves $\alpha$ and $\beta$. In his seminal 1988 Bulletin paper [Thu88], Thurston provides the upper bound of 6g66g-6 on the algebraic degree of a pseudo-Anosov stretch factor $\lambda(f)$ and claims, without proof, that “the examples of [Thu88, Theorem 7] show that this bound is sharp”. The referenced examples are exactly the pseudo-Anosov maps in $\langle T_{\alpha},T_{\beta}\rangle$.
Margalit remarked in 2011 what Strenner wrote down in his article on stretch factor degrees [Str17], namely that no proof of Thurston’s claim has ever been published. We are finally able to substantiate Thurston’s claim. Furthermore, we can even do so for pseudo-Anosov maps in the Torelli group. Our precise statement for the Thurston–Veech construction is the following.
Theorem 3.
Let $g\geq 2$ and $1\leq d\leq 3g-3$ be integers. Then there exists a pseudo-Anosov map on $S_{g}$ arising from the Thurston–Veech construction with trace field degree $d$ and stretch factor degree $2d$. For $g\geq 3$, the pseudo-Anosov maps can be chosen in the Torelli group $\mathcal{I}(S_{g})$.
Clearly, 3 implies 1 and 2 for $g\geq 3$. Our proof of the Torelli case of 3 does not work for $g=2$, and for this situation we directly prove 1 and 2 by using ad hoc examples (see Section 1.4).
1.3. Proof strategy for 3
For a pair $\alpha,\beta\subset S_{g}$ of filling multicurves, let $X=\left(|\alpha_{i}\cap\beta_{j}|\right)_{ij}$ be the matrix encoding the number of geometric intersections of the components of $\alpha$ and $\beta$.
The matrix $XX^{\top}$ is primitive, hence by Perron–Frobenius theory its spectral radius equals its largest eigenvalue and is therefore an algebraic integer. Let $d$ be its algebraic degree. We call the number $d$ the multicurve intersection degree of $\alpha$ and $\beta$.
Our proof is based on the following existence result.
Theorem 4.
Let $\alpha,\beta\subset S$ be a pair of filling multicurves having multicurve intersection degree $d$. For $\varepsilon\in\mathbb{Z}\setminus\{0\}$, there exists $n\in\mathbb{Z}_{>0}$ such that the mapping class $T^{n}_{\alpha}\circ T^{n\varepsilon}_{\beta}$ is pseudo-Anosov with stretch factor $\lambda$ of degree $2d$.
Assuming 4, what remains to be done in order to prove the first part of 3 is to construct all multicurve intersection degrees $1\leq d\leq 3g-3$ on $S_{g}$ for $g\geq 2$. By the Thurston–Veech construction, the trace field degree of the resulting pseudo-Anosov maps equals exactly the multicurve intersection degree of the multicurves $\alpha$ and $\beta$ used in the construction [Thu88, Vee89]. Hence, we are done by setting $\varepsilon=1$ in 4.
In order to prove the Torelli part of 3, we construct the multicurves $\alpha$ and $\beta$ realising the multicurve intersection degrees $1\leq d\leq 3g-3$ in such a way that $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of $\mathcal{I}(S_{g})$. We will ensure this by choosing multicurves $\alpha$ and $\beta$ which consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. We can then finish the proof of 3 by setting $\varepsilon=-1$ in 4. To see this, note that if $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of $\mathcal{I}(S_{g})$, then so is $T_{\alpha}^{n}\circ T_{\beta}^{-n}$.
We note that our examples of multicurves $\alpha$ and $\beta$ that we use in order to construct examples in the Torelli group $\mathcal{I}(S_{g})$ cannot yield multicurve intersection degrees greater than one on the surface $S_{2}$. Indeed, there exist no bounding pairs on $S_{2}$ and a multicurve can have at most one separating component.
The constructed examples in [Str17] do not belong to the Torelli group. Indeed, the first components of the considered multicurves $A$ and $B$ [Str17, Figure 6.1] are nonzero in homology and they intersect nontrivially.
To find the suitable multicurves $\alpha$ and $\beta$ realising all possible multicurve intersection degrees $1\leq d\leq 3g-3$ is the main technical contribution in this article.
1.4. Proof of 1 and 2 for $g=2$
For the genus two surface we give ad hoc examples. For this purpose, we use the flipper software [Bel21] and start with the genus two surface with one puncture $S_{2,1}$, see the figure below. In flipper, mapping classes are defined via Dehn twists along the curves $a,b,c,d,e,f$.
Consider the separating curve $\gamma$ depicted in blue. By the chain relation $T_{\gamma}=(T_{a}\circ T_{f})^{6}$. We consider the following three conjugates of $T_{\gamma}$:
  1. (1)
    $T_{1}=(T_{f}\circ T_{a}\circ T_{b})\circ T_{\gamma}\circ(T_{f}\circ T_{a}\circ T_{b})^{-1}$,
  2. (2)
    $T_{2}=(T_{c}\circ T_{b})\circ T_{\gamma}\circ(T_{c}\circ T_{b})^{-1}$,
  3. (3)
    $T_{3}=(T_{a}\circ T_{b})\circ T_{\gamma}\circ(T_{a}\circ T_{b})^{-1}$.
$a$
$b$
$c$
$f$
$d$
$e$
$\gamma$
Obviously $T_{\gamma},T_{1},T_{2},T_{3}\in\mathcal{I}(S_{2,1})$. For each even degree $d\in\{2,4,6\}$, we exhibit a word in the above elements. We check that the mapping class is pseudo-Anosov with singularity pattern $(1,1,1,1;0)$, and we compute its stretch factor by using flipper [Bel21]. The vector $(1,1,1,1;0)$ means that the invariant foliations have four $3$-pronged type singularities, and a $2$-pronged one at the puncture.
pseudo–Anosov mapping class $f]\in\mathrm{Mod}(S_{2,1})$ minimal polynomial of $\lambda(f)$
$T_{\gamma}\circ T_{1}\circ T_{2}^{-1}$ $t^{2}-66t+1$
$T_{\gamma}\circ T_{1}\circ T_{2}$ $t^{4}-72t^{3}+110t^{2}-72t+1$
$T_{\gamma}\circ T_{1}\circ T_{2}\circ T_{3}$ $t^{6}-266t^{5}+143t^{4}-204t^{3}+143t^{2}-266t+1$
In order to obtain elements in $\mathrm{Mod}(S_{2})$, we appeal to the forgetful map. Since our examples do not have a $1$-pronged singularity at the puncture, we can fill it in order to get a pseudo-Anosov mapping class, with the same stretch factor, in $\mathrm{Mod}(S_{2})$, see [HK06, Lemma 2.6] for details. This completes the proof of 2 for $g=2$. Now for each example, the minimal polynomial of $\lambda(f)$ is reciprocal and has degree $2d$ (for $d=1,2,3$ respectively). Hence the minimal polynomial of $\lambda(f)+\lambda(f)^{-1}$ has degree $d$ as required, and 1 is proved.
1.5. Explicit examples
In [Mar19, Problem 10.4], Margalit asks for explicit examples of pseudo-Anosov maps with specific stretch factor degrees. Our construction of multicurves allows us to do so for small genera.
More precisely, we first provide an almost explicit construction of multicurves with intersection degrees $1<d\leq 3g-3$ for which the pseudo-Anosov mapping class $T_{\alpha}\circ T_{\beta}$ has stretch factor degree $2d$. In the setting of 4, this means for $\varepsilon=1$ one can choose $n=1$. Our argument uses [LL24, Theorem 6] (see Section 6 for details). Unfortunately, for the multicurves we construct in the Torelli case, the criterion from [LL24] does not apply. Hence our need to use 4 for the proof of 3, which provides a slightly less explicit result.
Aided by the computer, we can subsequently find completely explicit examples of multicurves and therefore entirely explicit pseudo-Anosov mapping classes on $S_{g}$ arising from the Thurston–Veech construction realising the maximal stretch factor degree 6g66g-6, for all genera $g\leq 201$.
We emphasise that this construction, contained in Section 6, is independent of the proof of our main results, 1, 2 and 3.
1.6. Odd degree stretch factors.
While 4 provides the existence of field extensions $\mathbb{Q}(\lambda):\mathbb{Q}(\lambda+\lambda^{-1})$ of degree two for mapping classes in $\langle T_{\alpha},T_{\beta}\rangle$, realising extensions of degree one seems to be more mysterious. For example, Veech [Vee82] discovered a family of Hecke groups $\langle T_{\alpha},T_{\beta}\rangle=\langle\left(\begin{smallmatrix}1&\lambda_{q}\\ 0&1\end{smallmatrix}\right),\left(\begin{smallmatrix}1&0\\ -\lambda_{q}&1\end{smallmatrix}\right)\rangle$, where $\lambda_{q}=2\cos\pi/q$ for $q\geq 3$. The genus of the surface $S_{g}$ is (q1)/2(q-1)/2 for odd $q$. For $q=7,9$ one can find stretch factors of degree one over the trace field $\mathbb{Q}(\lambda_{q})$: for instance $T_{\alpha}\circ T_{\beta}^{-1}$ is an example for $q=7$, and we refer to [Bou22] for $q=9$. However, it is conjectured (see [HMTY08, Remark 9]) that stretch factors of degree one over $\mathbb{Q}(\lambda_{q})$ do not exist for odd $q\geq 11$.
It remains an open problem to construct odd stretch factor degrees in the Torelli group $\mathcal{I}(S_{g})$.
1.7. On Thurston’s upper bound
One may impose restrictions on the foliations fixed by the pseudo-Anosov map, for example by prescribing the stratum, that is, the number of singularities as well as their orders as $k$-pronged singularities. It turns out that Thurston’s upper bound for the stretch factor degree becomes
$2g-2+\#\{\textrm{odd singularities}\}\leq 2g-2+4g-4=6g-6.$
In the context of the Thurston–Veech construction, the number and type of singularities can be read off directly from the geometry of the complement $S\setminus(\alpha\cup\beta)$: the number of $k$-pronged singularities coincides with the number of $2k$-gons in the complement.
It is a consequence of our proof of 3 that the upper bound
$2g-2+\#\{\textrm{odd singularities}\}$
for the stretch factor degree is sharp for every $g\geq 3$, and that examples realising this upper bound can be taken in the Torelli group.
1.8. A natural field extension from the perspective of curves
A pair of filling multicurves $\alpha,\beta\subset S_{g}$ naturally determines a bipartite graph whose vertices correspond to curve components and the number of edges between each pair of vertices equals the number of intersection points of the respective curve components. The adjacency matrix of this graph is $\Omega=\left(\begin{smallmatrix}0&X\\ X^{\top}&0\end{smallmatrix}\right)$. Clearly, the square root $\sqrt{\mu}$ of the spectral radius $\mu$ of $XX^{\top}$ equals the spectral radius of $\Omega$. We call the algebraic degree of $\sqrt{\mu}$ the multicurve bipartite degree of $\alpha$ and $\beta$. Similarly to the field extension $\mathbb{Q}(\lambda):\mathbb{Q}(\lambda+\lambda^{-1})$, also the field extension $\mathbb{Q}(\sqrt{\mu}):\mathbb{Q}(\mu)$ has degree one or two. From the point of view of the Thurston–Veech construction one has $\lambda(f)+\lambda(f)^{-1}=2+\mu$ for $f=T_{\alpha}\circ T_{\beta}^{-1}$ (see the proof of 4). We prove the following result.
Theorem 5.
Every even integer $2\leq 2d\leq 6g-6$ is realised as a multicurve bipartite degree on $S_{g}$ for $g\geq 2$.
Organisation
In Section 2 we prove 4, which is the nonsplitting criterion used to reduce 3 to the construction of certain kinds of pairs of multicurves. In Section 3 we introduce an irreducibility criterion for the characteristic polynomial of matrices of the form $XX^{\top}$ which plays a central role throughout the rest of the article. Using this irreducibility criterion, we first give a proof of Thurston’s claim as a warm-up in Section 4, before providing the multicurves needed to prove 3 in Section 5. Finally, we provide some explicit examples in Section 6.
Acknowledgments
The authors thank Dan Margalit and Jean-Claude Picaud for inspiring discussions, Mark Bell for his help with the software flipper [Bel21] and Curt McMullen for comments on an earlier version of the article. They are particularly grateful to Dan Margalit for pointing out a mistake in the formulation of the proof of 3 in an earlier version of the article. The authors also thank the anonymous referee for several useful suggestions, comments and corrections. The first author has been partially supported by the LabEx PERSYVAL-Lab (ANR-11-LABX-0025-01) funded by the French program Investissement d’avenir. The first author would like to thank the University of Fribourg and the Center for Mathematical Modeling (CMM) at Universidad de Chile for excellent working conditions.
2. A nonsplitting criterion
In this section we prove 4, which is an algebraic criterion that allows us to deduce that the degree of the field extension $\mathbb{Q}(\lambda(f)):\mathbb{Q}(\lambda(f)+\lambda(f)^{-1})$ equals two for certain $f$ which are a product of multitwists. Compare with [LL24, Theorem 6]. For convenience, we repeat the statement of 4:
Theorem (4).
Let $\alpha,\beta\subset S$ be a pair of filling multicurves having multicurve intersection degree $d$. For every $\varepsilon\in\mathbb{Z}\setminus\{0\}$, there exists $n\in\mathbb{Z}_{>0}$ such that the mapping class $T^{n}_{\alpha}\circ T^{n\varepsilon}_{\beta}$ is pseudo-Anosov with stretch factor $\lambda$ of degree $2d$.
We need the following lemma which is a small extension of a result of Murty [Mur08, Theorem 4]. It is probably well-known but we could not find a reference so we include a proof here.
Lemma 6.
Let $Q(t)\in\mathbb{Z}[t]$ be a polynomial taking an integral square value at all but finitely many positive integer specialisations. Then $Q(t)$ is the square of a polynomial in $\mathbb{Z}[t]$.
Proof.
Let $N\in\mathbb{N}$ be a positive integer such that $Q(n)$ is a square for every $n\in\mathbb{N}$ with $n>N$. Let $n\in\mathbb{Z}$. Consider a prime number $p$ such that $n+p>N$. The integer $Q(n+p)$ is a square by assumption, and its reduction modulo $p$ is $Q(n)$. Thus, $Q(n)\in\mathbb{Z}$ is a quadratic residue for all but finitely many prime numbers. So the value $Q(n)$ is a square in $\mathbb{Z}$ (see [Hal33, Theorem 1 (end of the proof)]). Thus we can apply [Mur08, Theorem 4] to conclude that $Q(t)$ is the square of a polynomial in $\mathbb{Z}[t]$. ∎
Proof of 4.
By the Thurston–Veech construction, there exists a representation $\rho:\left<T_{\alpha},T_{\beta}\right>\to\mathrm{PSL}_{2}(\mathbb{R})$ mapping $T_{\alpha}$ to the matrix $\left(\begin{smallmatrix}1&r\\ 0&1\end{smallmatrix}\right)$ and $T_{\beta}$ to the matrix $\left(\begin{smallmatrix}1&0\\ -r&1\end{smallmatrix}\right)$, where $r^{2}=\mu$ is the spectral radius of the matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$. Furthermore, the stretch factor $\lambda(f)$ of $f\in\left<T_{\alpha},T_{\beta}\right$> equals the spectral radius of $\rho(f)$. Now, let us consider the product of multitwists $f=T^{2n}_{\alpha}\circ T^{2n\varepsilon}_{\beta}$. A direct computation provides that the trace of $\rho(f)$ equals $\mathrm{tr}(\rho(f))=2-\varepsilon(2nr)^{2}$. Thus, $\lambda(f)+\lambda(f)^{-1}=|2-\varepsilon(2nr)^{2}|$ and hence $\mathbb{Q}(\lambda(f)+\lambda(f)^{-1})=\mathbb{Q}(\mu)=K$. Note that by assumption, the degree of the field extension $K:\mathbb{Q}$ is $d$, the multicurve intersection degree of $\alpha$ and $\beta$.
Since $\lambda=\lambda(f)$ solves the quadratic equation $t^{2}-(\lambda+\lambda^{-1})t+1=0$$\lambda$ has degree $1$ or $2$ over $K$. All that we need to do is find $n\in\mathbb{Z}_{>0}$ such that $\lambda\not\in K$, or equivalently such that the discriminant $D=(2-\varepsilon(2nr)^{2})^{2}-4=16\cdot n^{2}\cdot((n\varepsilon\mu)^{2}-\varepsilon\mu)$ of the quadratic equation is not a square in $K$. We proceed by contradiction. Let $\mu^{\prime}=\varepsilon\mu$ and let us assume that $(n\mu^{\prime})^{2}-\mu^{\prime}$ is a square in $K=\mathbb{Q}(\mu^{\prime})$ for every $n>0$. Since the expression $(n\mu^{\prime})^{2}-\mu^{\prime}$ is invariant under the transformation $n\mapsto-n$, we can assume that $(n\mu^{\prime})^{2}-\mu^{\prime}$ is a square for every $n\in\mathbb{Z}\setminus\{0\}$.
Let $q\in\mathbb{Q}[t]$ be the minimal polynomial of $\frac{1}{\mu^{\prime}}$ over $\mathbb{Q}$. Our strategy is to use 6 to show the following claim.
Claim. The polynomial $q(t^{2})\in\mathbb{Q}[t]$ has a double root in $\mathbb{C}$.
This claim directly leads to a contradiction. Indeed, the polynomial $q\in\mathbb{Q}[t]$ is irreducible, and thus only has simple roots. Now each root $0\neq a\in\mathbb{C}$ of $q$ gives rise to two distinct roots $\pm\sqrt{a}$ of $q(t^{2})$, contradicting the claim. In order to finish the proof of 4, it therefore remains to prove the claim.
Proof of Claim. Let $P=a_{d}t^{d}+a_{d-1}t^{d-1}+\dots+a_{1}t+a_{0}\in\mathbb{Q}[t]$ be the minimal polynomial of $\mu^{\prime}$ over $\mathbb{Q}$ with $a_{d}=1$. The Thurston–Veech construction implies that $\mu$ is an eigenvalue of a square matrix with integer coefficients, so $\mu^{\prime}$ is also an eigenvalue of a square matrix with integer coefficients, implying that $P\in\mathbb{Z}[t]$.
Thus, $\mu^{\prime}$ and $n^{2}\mu^{\prime}-1$ are algebraic integers, and we can consider their norms $N(\cdot)$. Obviously, for $\mu^{\prime}$ one has $N(\mu^{\prime})=(-1)^{d}a_{0}$. Let us compute the norm of $n^{2}\mu^{\prime}-1$. The monic polynomial $n^{2d}P\left(\frac{t+1}{n^{2}}\right)\in\mathbb{Z}[t]$ has degree $d$ and admits $n^{2}\mu^{\prime}-1$ as a root. Moreover the fields $\mathbb{Q}(\mu^{\prime})$ and $\mathbb{Q}(n^{2}\mu^{\prime}-1)$ coincide, so they have the same degree over $\mathbb{Q}$, namely $d$. Hence $n^{2d}P\left(\frac{t+1}{n^{2}}\right)$ is the minimal polynomial of $n^{2}\mu^{\prime}-1$. Inspecting its constant term, we deduce
$N(n^{2}\mu^{\prime}-1)=(-1)^{d}\sum_{k=0}^{d}a_{k}n^{2d-2k}.$
Since the norm is multiplicative, we have
$N((n\mu^{\prime})^{2}-\mu^{\prime})=N(n^{2}\mu^{\prime}-1)\cdot N(\mu^{\prime})=a_{0}\sum_{k=0}^{d}a_{k}n^{2d-2k}=Q(n^{2}),$
where
$Q(t)=a_{0}\sum_{k=0}^{d}a_{k}t^{d-k}.$
Since by assumption $(n\mu^{\prime})^{2}-\mu^{\prime}$ is a square for every $n\in\mathbb{Z}\setminus\{0\}$, we deduce that $Q(n^{2})$ is a square for every $n\in\mathbb{Z}\setminus\{0\}$. By 6, $Q(t^{2})$ is itself a square in $\mathbb{Z}[t]$.
To finish the proof, observe that $Q(t)=a_{0}\cdot t^{d}P\left(\frac{1}{t}\right)\in{\mathbb{Z}}[t]$. In particular, $Q\left(\frac{1}{\mu^{\prime}}\right)=0$ and $Q$ has degree $d$ over $\mathbb{Q}$. Since $\mu^{\prime}$ and $\frac{1}{\mu^{\prime}}$ generate the same field extension $K:\mathbb{Q}$, the minimal polynomial $q$ of $\frac{1}{\mu^{\prime}}$ over $\mathbb{Q}$ is also of degree $d$. This implies that $Q$ must be a rational multiple of $q$. In particular, $q(t^{2})$ is a rational multiple of $Q(t^{2})$ and therefore has a double root in $\mathbb{C}$. ∎
3. An irreducibility criterion
The goal of this section is to present an algebraic criterion that allows us to deduce that certain characteristic polynomials of matrices of the form $XX^{\top}$ are irreducible. We write $E_{11}$ for the matrix with the top-left coefficient equal to 11 and all other coefficients zero. Furthermore, for a square matrix $A$ with coefficients in a ring $R$, we denote its characteristic polynomial by $\chi_{A}=\det(t\cdot\mathrm{Id}-A)\in R[t]$.
Proposition 7.
Let $M$ be a square integer matrix, and let $N$ be the principal submatrix of $M$ obtained by deleting the first row and the first column. If $M$ and $N$ have no common eigenvalue, and if $M$ has a simple eigenvalue $\rho$, then the characteristic polynomial of $\widetilde{M}=M+{c}y^{p}E_{11}$ is an irreducible element of $\mathbb{Z}[t,y]$, for all $0\neq p\in\mathbb{N}$ and $0\neq{c}\in\mathbb{Z}$.
Proof.
Our goal is to use Eisenstein’s criterion on $\chi_{\widetilde{M}}\in\mathbb{Z}[t,y]\cong\left(\mathbb{Z}[t]\right)[y]$, viewing it as a polynomial in the variable $y$ and coefficients in $\mathbb{Z}[t]$. By multilinearity of the determinant, we calculate
$\chi_{\widetilde{M}}(t,y)=\det(t\cdot\mathrm{Id}-{\widetilde{M}})=-y^{p}{c}\chi_{N}(t)+\chi_{M}(t)$
and notice that ${c}\chi_{N}$ and $\chi_{M}$ are relatively prime in $\mathbb{Z}[t]$. Indeed, $\chi_{M}$ has leading coefficient +1+1 and no root in common with $\chi_{N}$ by our assumption that $M$ and $N$ have no eigenvalue in common. In particular, they have no common factor, which shows that $\chi_{\widetilde{M}}\in\left(\mathbb{Z}[t]\right)[y]$ is primitive. In order to apply Eisenstein’s criterion, let $\mu_{\rho}\in\mathbb{Z}[t]$ be the minimal polynomial of the simple eigenvalue $\rho$ of $M$. By assumption, $\mu_{\rho}$ divides $\chi_{M}$ exactly once, but it does not divide $\chi_{N}$ since $\chi_{M}$ and $\chi_{N}$ have no common root. In particular, Eisenstein’s criterion applies to show that the polynomial $\chi_{\widetilde{M}}\in\left(\mathbb{Z}[t]\right)[y]\cong\mathbb{Z}[t,y]$ is irreducible. ∎
Remark 8.
In the previous statement, one can easily replace $\chi_{\widetilde{M}}(t)$ by $\chi_{\widetilde{M}}(t^{n})$ for any integer $n>0$. Indeed $\chi_{M}(t^{n})$ and $\chi_{N}(t^{n})$ are still coprime and $\mu_{\rho}(t^{n})$ divides $\chi_{M}(t^{n})$ exactly once, so Eisenstein’s criterion applies.
Remark 9.
Oscillatory matrices satisfy a stronger version of Perron–Frobenius theory, namely all the eigenvalues are positive real, simple, and they strictly interlace when taking a principal submatrix [And87]. Hence, 7 applies very cleanly to this class of matrices.
We use 7 on the following two cases (10 and 11).
Lemma 10.
For $n\geq 1$, let
$N=\left(\begin{array}[]{c|ccc}a_{1}&a_{2}&\dots&a_{n}\\ \hline\cr a_{2}&&&\\ \vdots&&\ast&\\ a_{n}&&&\end{array}\right),\hskip 14.22636ptM=\left(\begin{array}[]{c|ccc}0&\alpha a_{1}&\dots&\alpha a_{n}\\ \hline\cr\alpha a_{1}&&&\\ \vdots&&N&\\ \alpha a_{n}&&&\end{array}\right)$
be square integer matrices with $a_{1}\geq 1$ and $0\neq\alpha\in\mathbb{Q}$. If $M$ is nonnegative and irreducible, and if $\chi_{N}\in\mathbb{Z}[t]$ is irreducible, then the characteristic polynomial of $\widetilde{M}=M+{c}y^{2}E_{11}$ is irreducible in $\mathbb{Z}[t,y]$ for all $0\neq{c}\in\mathbb{Z}$.
Proof.
In order to use 7, we need to show that $M$ has a simple eigenvalue and that $M$ and $N$ share no eigenvalue. The former holds since $M$ is nonnegative and irreducible, and in particular has a Perron–Frobenius eigenvalue which is simple, see for example [Gan59]. For the latter, we compute
$\chi_{M}(t)=t\chi_{N}(t)+q(t),$
where $q(t)\in\mathbb{Z}[t]$ is of degree at most n1n-1. We claim that it is not the zero polynomial either. Indeed, we directly verify
$q(0)$ $=\det\left(\begin{array}[]{c|ccc}0&-\alpha a_{1}&\dots&-\alpha a_{n}\\ \hline\cr-\alpha a_{1}&&&\\ \vdots&&-N&\\ -\alpha a_{n}&&&\end{array}\right)$
$=\det\left(\begin{array}[]{c|ccc}\alpha^{2}a_{1}&0&\dots&0\\ \hline\cr 0&&&\\ \vdots&&-N&\\ 0&&&\end{array}\right)=\pm\alpha^{2}a_{1}\det N\neq 0.$
Now if there existed a common root $\lambda\in\mathbb{C}$ of $\chi_{M}$ and $\chi_{N}$, then $\lambda$ would also be a root of $q(t)$, as the equation $\chi_{M}(t)=t\chi_{N}(t)+q(t)$ implies $q(\lambda)=\chi_{M}(\lambda)-\lambda\chi_{N}(\lambda)=0$. Since $\chi_{N}$ is irreducible, it is the minimal polynomial of $\lambda$ over $\mathbb{Q}$, and we assume it to be of degree $n$. But $q(t)\in\mathbb{Z}[t]$ is a nonzero polynomial of degree at most n1n-1 with root $\lambda$, which is impossible by the definition of the minimal polynomial. Hence, $\chi_{M}$ and $\chi_{N}$ have no common root $\lambda\in\mathbb{C}$. ∎
Lemma 11.
For $n,m\geq 1$, let
$A=\left(\begin{array}[]{c|ccc}a_{1}&a_{2}&\dots&a_{n}\\ \hline\cr a_{2}&&&\\ \vdots&&\ast&\\ a_{n}&&&\end{array}\right),\hskip 14.22636ptB=\left(\begin{array}[]{c|ccc}b_{1}&b_{2}&\dots&b_{m}\\ \hline\cr b_{2}&&&\\ \vdots&&\ast&\\ b_{m}&&&\end{array}\right)$
be square integer matrices of size $n$ and $m$, respectively, with $a_{1},b_{1}\geq 1$. Furthermore, let $\alpha,\beta\neq 0$ such that
$M=\left(\begin{array}[]{c|ccc|ccc}0&\alpha a_{1}&\dots&\alpha a_{n}&\beta b_{1}&\dots&\beta b_{m}\\ \hline\cr\alpha a_{1}&&&&&&\\ \vdots&&A&&&&\\ \alpha a_{n}&&&&&&\\ \hline\cr\beta b_{1}&&&&&&\\ \vdots&&&&&B&\\ \beta b_{m}&&&&&&\end{array}\right)$
is a matrix with integer coefficients. If $M$ is nonnegative and irreducible, and if $\chi_{A},\chi_{B}\in\mathbb{Z}[t]$ are irreducible and distinct, then the characteristic polynomial of $\widetilde{M}=M+{c}y^{2}E_{11}$ is irreducible in $\mathbb{Z}[t,y]$ for all $0\neq{c}\in\mathbb{Z}$.
Proof.
The proof is similar to the proof of 10: the only thing to verify is that no eigenvalue of $A$ or of $B$ is also an eigenvalue of $M$. Again, we compute
$\chi_{M}(t)=t\chi_{A}(t)\chi_{B}(t)\pm q_{1}(t)\chi_{B}(t)\pm q_{2}(t)\chi_{A}(t),$
where $q_{1}(t)\in\mathbb{Z}[t]$ is of degree at most n1n-1 and $q_{2}(t)\in\mathbb{Z}[t]$ is of degree at most m1m-1. This is seen by developing the first column of the matrix tIMtI-M. The first coefficient is responsible for the summand $t\chi_{A}(t)\chi_{B}(t)$, the next $n$ coefficients are responsible for the summand $\pm q_{1}(t)\chi_{B}(t)$ and the final $m$ coefficients are responsible for the summand $\pm q_{2}(t)\chi_{A}(t)$. We claim that neither among $q_{1}(t)$ and $q_{2}(t)$ is the zero polynomial. Indeed, by developing the first column of the matrix $t\mathrm{I}-M$, and evaluating at $t=0$, we get
$q_{1}(0)$ $=\det\left(\begin{array}[]{c|ccc}0&-\alpha a_{1}&\dots&-\alpha a_{n}\\ \hline\cr-\alpha a_{1}&&&\\ \vdots&&-A&\\ -\alpha a_{n}&&&\end{array}\right)$
$=\det\left(\begin{array}[]{c|ccc}\alpha^{2}a_{1}&0&\dots&0\\ \hline\cr 0&&&\\ \vdots&&-A&\\ 0&&&\end{array}\right)=\pm\alpha^{2}a_{1}\det A,$
which is not zero since $\chi_{A}$ is irreducible. Similarly, $q_{2}(0)\neq 0$. Now if there existed a common root $\lambda\in\mathbb{C}$ of $\chi_{M}$ and $\chi_{A}$, then $\lambda$ would also be a root of either $\chi_{B}$ or $q_{1}$. Since $\chi_{A}$ and $\chi_{B}$ are irreducible and distinct, we must have $q_{1}(\lambda)=0$. But since $\chi_{A}$ is irreducible of degree $n$, and $q_{1}(t)$ is a nonzero polynomial of degree at most n1n-1, this is impossible. Similarly, no root of $\chi_{B}$ can be a root of $\chi_{M}$, which concludes the proof. ∎
Remark 12.
One could formulate 11 with $k\geq 2$ blocks $A_{1},\dots,A_{k}$ of respective sizes $n_{1},\dots,n_{k}$, instead of $k=2$. In this case, all the $k$ characteristic polynomials $\chi_{A_{i}}$ should be irreducible and pairwise distinct. The argument is identical by considering χ_ M (t) = t ∏_i=1^k χ_A_i + ∑_i=1^k ±q_i(t) ∏_j≠i χ_A_j, where $q_{i}(t)\in\mathbb{Z}[t]$ is of degree at most $n_{i}-1$ and nonzero.
4. Warm-up: proof of Thurston’s claim
As a first illustration of our method, it is the goal of this section to provide a pair of filling multicurves $\alpha$ and $\beta$ on $S_{g}$ with multicurve intersection degree 3g33g-3. By 4, this validates Thurston’s claim that the product of two multitwists can realise the maximal possible algebraic degrees of stretch factors: 6g66g-6.
Recall that the matrix encoding the number of intersections of the components of $\alpha$ and $\beta$ is denoted by $X=\left(|\alpha_{i}\cap\beta_{j}|\right)_{ij}$ (see Section 1.3). In order to read off the matrix $XX^{\top}$ from our figures, we use the following formula for its coefficients, which is a direct consequence of the definition of matrix multiplication:
$(XX^{\top})_{ij}=\sum_{k}|\alpha_{i}\cap\beta_{k}|\cdot|\beta_{k}\cap\alpha_{j}|.$
We start by realising, on the surface of genus $g\geq 1$ with $2$ boundary components, a pair of filling multicurves $\alpha$ and $\beta$ such that $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree 3g13g-1. On a surface with boundary, we say that a pair of multicurves is filling if they intersect transversally and if their complement consists of boundary-parallel annuli and of discs none of which is a bigon. We proceed by induction on $g$.
For $g=1$ with two boundary components
We consider the filling pair of multicurves $\alpha$ and $\beta$ shown in Figure 1, where one of the components of $\beta$ has y1y-1 parallel copies. Here, we think of $y$ as a variable that we specify to a value in $\mathbb{Z}_{\geq 2}$ later on.
y1y-1 copies
$\alpha_{2}$
$\alpha_{1}$
Figure 1. Two multicurves $\alpha$ (in red) and $\beta$ (in blue) on the surface of genus one with two boundary components. The multicurve $\beta$ contains y1y-1 parallel copies of one of its components.
Observe that $X$ is a matrix of size $2\times y$ (the multicurve $\beta$ has $y$ components). We directly calculate
$XX^{\top}=\begin{pmatrix}4&2\\ 2&y\end{pmatrix}$.
We have $\chi_{XX^{\top}}(t)=t^{2}-(4+y)t+4(y-1)$ with discriminant $y^{2}-8y+32$, which is not a square if $y\geq 12$. Indeed, in this case we have
$(y-3)^{2}=y^{2}-6y+9>y^{2}-8y+32>y^{2}-8y+16=(y-4)^{2}.$
In particular, for $y\geq 12$ the polynomial $\chi_{XX^{\top}}$ is irreducible.
For $g>1$ and two boundary components
For the inductive step, assume we have constructed on the surface of genus $g\geq 1$ with $2$ boundary components a filling pair of multicurves $\alpha^{\prime},\beta^{\prime}$ such that the characteristic polynomial $\chi^{\prime}=\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree 3g13g-1. Furthermore, we inductively assume that $\alpha^{\prime}_{1}$ is a simple closed curve that encircles all the handles of the surface, as illustrated in Figure 2. Take a surface of genus 11 and two boundary components, as in the case of genus $g=1$, see Figure 1, and denote its filling pair of multicurves by $\alpha^{\prime\prime}$ and $\beta^{\prime\prime}$. We assume that the component of $\beta^{\prime\prime}$ with multiple parallel copies has b1b-1 of them. Now glue its right boundary component to the left boundary component of the surface of genus $g$, and add two new curves $\alpha_{0}$ and $\beta_{0}$ to the multicurves. The curve $\alpha_{0}$ encircles all the handles of the newly formed surface, and the curve $\beta_{0}$ twice intersects $\alpha_{0}$ but no other multicurve component. Again, see Figure 2 for an illustration. By construction, the pair of multicurves $\alpha,\beta$ is again filling.
$\cdots$
$\beta_{0}$
$\alpha_{0}$
$\alpha_{1}^{\prime}$
$\alpha_{1}^{\prime\prime}$
Figure 2. Two surfaces of genus 11 and $g$, respectively, and two boundary components, glued together along one of their boundary components. The curves $\alpha^{\prime}_{1}$ and $\alpha^{\prime\prime}_{1}$ are shown, each encircling all the handles of their respective surface. The new curve $\alpha_{0}$ encircles all the handles of the newly formed surface, and the new curve $\beta_{0}$ runs along the glued boundary component.
Let $A$ be the matrix $XX^{\top}$ for the pair of multicurves $\alpha^{\prime},\beta^{\prime}$, and let $B$ be the matrix $XX^{\top}$ for the pair of multicurves $\alpha^{\prime\prime},\beta^{\prime\prime}$. We define the multicurves
$\alpha$ $=\alpha_{0}\cup\alpha^{\prime}\cup\alpha^{\prime\prime},$
$\beta$ $=\beta_{0}\cup\beta^{\prime}\cup\beta^{\prime\prime}.$
A quick computation gives
$A=\left(\begin{array}[]{c|ccc}a_{1}&a_{2}&\dots&a_{n}\\ \hline\cr a_{2}&&&\\ \vdots&&\ast&\\ a_{n}&&&\end{array}\right),\hskip 14.22636ptB=\left(\begin{array}[]{cccc}4&2\\ 2&b\end{array}\right).$
Let us choose $b$ such that $\chi_{B}$ is irreducible and distinct from $\chi_{A}$. We note $a_{1}=4a$ for a positive integer $a$. Indeed, since $\alpha^{\prime}_{1}$ is separating, its geometric intersection number with each component $\beta^{\prime}_{k}$ is a multiple of two, say $2x_{k}$. In particular, we have
$a_{1}=\sum_{k}|\alpha^{\prime}_{1}\cap\beta^{\prime}_{k}|^{2}=4\sum_{k}x_{k}=4a.$
In the multicurve $\beta$, we take $y^{2}-a-1\geq 1$ parallel copies of $\beta_{0}$, for $y>0$ large enough. The matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$ takes the form
$XX^{\top}=\left(\begin{array}[]{c|ccc|cc}4y^{2}&a_{1}&\dots&a_{n}&4&2\\ \hline\cr a_{1}&&&&&\\ \vdots&&A&&&\\ a_{n}&&&&&\\ \hline\cr 4&&&&4&2\\ 2&&&&2&b\end{array}\right).$
It is a crucial feature of the Thurston–Veech construction that for a filling pair of multicurves, such a matrix $XX^{\top}$ is always primitive and hence irreducible [Thu88, Vee89]. By 11, the polynomial $\chi_{XX^{\top}}\in\mathbb{Z}[t,y]$ is irreducible (recall that $\chi_{A}$ and $\chi_{B}$ are irreducible). Hence, by Hilbert’s irreducibility theorem, there exist infinitely many specifications of $y$ (and in particular infinitely many specifications of $y$ such that $y^{2}-a-1>0$) with $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ irreducible. The degree of this polynomial equals the size of the matrix $XX^{\top}$, which equals
3g1+3=3(g+1)1,3g-1+3=3(g+1)-1,
which is exactly what we wanted to show. Finally, to justify our inductive assumption on the curve $\alpha^{\prime}_{1}$, note that $\alpha_{0}$ is again a curve that encircles all handles of the newly built surface of genus $g+1$ with two boundary components.
The closed case for $g\geq 2$.
Take any example of a filling pair of multicurves $\alpha^{\prime}$ and $\beta^{\prime}$ we constructed on the surface of genus $g-1\geq 1$ with two boundary components. In particular, the curve $\alpha^{\prime}_{1}$ encircles all the g1g-1 handles of the surface, and if
$A=\left(\begin{array}[]{c|ccc}a_{1}&a_{2}&\dots&a_{n}\\ \hline\cr a_{2}&&&\\ \vdots&&\ast&\\ a_{n}&&&\end{array}\right)$
is the matrix $XX^{\top}$ for the multicurves $\alpha^{\prime}$ and $\beta^{\prime}$, we have $a_{1}=4a$ and $\chi_{A}(t)$ is irreducible. We identify the two boundary components of the surface to increase the genus by one. Let $\alpha_{0}$ be a longitude of the created handle, and let $\beta_{0}$ run along the glued boundary, see Figure 3. Define the two new multicurves
$\alpha$ $=\alpha_{0}\cup\alpha^{\prime}$
$\beta$ $=\beta_{0}\cup\beta^{\prime},$
where we take $y^{2}-a$ copies of $\beta_{0}$.
$\cdots$
$\beta^{\prime}_{1}$
$\alpha^{\prime}_{1}$
$\beta_{0}$
$\beta_{0}$
$\alpha_{0}$
Figure 3. The closed surface of genus $g$ is obtained by identifying the left and the right boundary component of the depicted closed surface of genus g1g-1 with two boundary components. The curve $\alpha_{0}$ is a longitude of the created handle, and the curve $\beta_{0}$ runs along the glued boundary.
By construction, the pair of multicurves $\alpha,\beta$ is again filling. The matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$ takes the form
$XX^{\top}=\left(\begin{array}[]{c|ccc}y^{2}&\frac{a_{1}}{2}&\dots&\frac{a_{n}}{2}\\ \hline\cr\frac{a_{1}}{2}&&&\\ \vdots&&A&\\ \frac{a_{n}}{2}&&&\end{array}\right),$
and $\chi_{XX^{\top}}\in\mathbb{Z}[t,y]$ is irreducible by 10. By Hilbert’s irreducibility theorem, there exist infinitely many specifications of $y$ (and in particular infinitely many specifications of $y$ such that $y^{2}-a>0$) with $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ irreducible. The degree of this polynomial equals the size of the matrix $XX^{\top}$, which equals
3(g1)1+1=3g3.3(g-1)-1+1=3g-3.
5. Proof of 3
The goal of this section is to realise every positive integer $d\leq 3g-3$ as the multicurve intersection degree of a pair of multicurves $\alpha,\beta\subset S$ on $S_{g}$ for $g\geq 3$ in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. This guarantees that $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of the Torelli group $\mathcal{I}(S_{g})$. Indeed, on the homology level, the action of a Dehn twist along a curve $\gamma$ is given by $\delta]\mapsto[\delta]+i(\delta,\gamma)[\gamma]$, where $i(\cdot,\cdot)$ is the algebraic intersection form on ${H}_{1}(S_{g};\mathbb{Z})$. It follows that a Dehn twist along a separating curve is en element of $\mathcal{I}(S_{g})$, but also each composition of Dehn twists $T_{\alpha_{i}}\circ T_{\beta_{j}}^{-1}$ for a bounding pair $\alpha_{i}$ and $\beta_{j}$. Since all the individual Dehn twists $T_{\alpha_{i}}$ commute, as well as all the individual Dehn twists $T_{\beta_{j}}$, it follows that if the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$, the mapping class $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of $\mathcal{I}(S_{g})$, and hence so is the mapping class $T_{\alpha}^{n}\circ T_{\beta}^{-n}$. 4 then provides 3 in the case $g\geq 3$. For the case $g=2$ we note that the statement is proved for d=3g3=3d=3g-3=3 in Section 4, and for $d\leq 2=g$ it is proved in [LL24].
We start with the maximal degree 3g33g-3 and then discuss how to adapt the construction in order to realise smaller degrees.
5.1. Multicurve intersection degree 3g33g-3
We start by realising, on the surface of genus $g\geq 2$ with one boundary component, a filling pair of multicurves $\alpha$ and $\beta$ such that $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree 3g23g-2, in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. The construction is done by induction on the genus $g\geq 2$.
5.1.1. For $g=2$ with one boundary component.
We consider the two multicurves $\alpha$ and $\beta$ shown in Figure 4, with $y$ parallel copies of the component $\beta_{1}$.
$\alpha_{1}$
$\alpha_{2}$
$\alpha_{3}$
$\alpha_{4}$
$y$ copies
of $\beta_{1}$
$\beta_{2}$
$\beta_{4}$
$\beta_{3}$
Figure 4. Two multicurves $\alpha$ and $\beta$ on the surface of genus two with one boundary component. There are $y$ parallel copies of the component $\beta_{1}$, and the separating component $\beta_{3}$ is drawn in light blue.
We first note that the components $\alpha_{1},\beta_{1},\alpha_{3}$ and $\beta_{3}$ are separating. Furthermore, the components $\alpha_{2}$ and $\alpha_{4}$ have their counterparts $\beta_{2}$ and $\beta_{4}$, respectively, in the multicurve $\beta$ with which they each form a bounding pair.
We directly calculate
$XX^{\top}=\begin{pmatrix}84+16y&40+8y&40&16\\ 40+8y&20+4y&20&8\\ 40&20&20&8\\ 16&8&8&4\end{pmatrix}$,
and it is a direct check (by the computer) that the characteristic polynomial of $XX^{\top}$ is irreducible if $y=2$ or $y=3$. This finishes the case $g=2$ with one boundary component.
5.1.2. For $g>2$ and one boundary component
In order to increase the genus by one, we glue a surface of genus one with two boundary components as follows. On this surface, we consider the two multicurves $\alpha$ and $\beta$ shown in Figure 5.
$\alpha_{1}$
$\alpha_{2}$
$y$ copies
Figure 5. Two multicurves $\alpha$ (in red) and $\beta$ (in blue) on the surface of genus one with two boundary components. The multicurve $\beta$ has $y$ parallel copies of its separating component.
We directly calculate
$XX^{\top}=\begin{pmatrix}16y+4&8y\\ 8y&4y\end{pmatrix}=:C_{y},$
and $\chi_{XX^{\top}}(t)=t^{2}-(20y+4)t+16y$ with discriminant $16\cdot(25y^{2}+6y+1)$, which is never a square. Indeed, we have
$(5y)^{2}=25y^{2}<25y^{2}+6y+1<25y^{2}+10y+1=(5y+1)^{2}.$
In particular, the polynomial $\chi_{XX^{\top}}$ is irreducible for every positive integer $y$.
For the inductive step, let $g\geq 2$. Assume we have constructed on the surface of genus $g$ with one boundary component a pair of multicurves $\alpha^{\prime},\beta^{\prime}$ such that the characteristic polynomial $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree 3g23g-2, in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. Further, assume that $\alpha_{1}^{\prime}$ is a simple closed curve that encircles all the handles of the surface, except for the rightmost one. Then, we take such a model surface and glue to its boundary a surface of genus one with two boundary components, as shown in Figure 5, and add two new curves $\alpha_{0}$ and $\beta_{0}$ to the multicurves. The curve $\alpha_{0}$ encircles all the handles of the newly formed surface, except for the rightmost one, and the curve $\beta_{0}$ runs along the glued boundary components, and twice intersects $\alpha_{0}$ but no other component of $\alpha$, see Figure 6.
$\cdots$
$\alpha_{1}^{\prime}$
$\alpha_{0}$
$\beta_{0}$
$\alpha^{\prime\prime}_{1}$
Figure 6.
The proof of irreducibility is now exactly the same as in the non-Torelli case in Section 4. The only thing we need to check is that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. But this is clearly the case, since all the curves we add in the inductive step are separating or come as a bounding pair.
5.1.3. The closed case for $g\geq 4$.
The last step is to make the surfaces closed. We simply glue together two pieces of genera $g^{\prime},g^{\prime\prime}$, where $g^{\prime}+g^{\prime\prime}=g$, and one boundary component together along their boundaries. The same argument as in the inductive step provides irreducible characteristic polynomials of degree
$3g^{\prime}-2+3g^{\prime\prime}-2+1=3g-3.$
5.1.4. The closed case for $g=3$.
We need a different argument. In this case, we start with the surface of genus two and one boundary component depicted in Figure 4, and close it off to the left by glueing a surface of genus one with one boundary component, see Figure 7.
$\alpha_{1}$
$\alpha_{2}$
$\alpha_{3}$
$\alpha_{4}$
22 copies
of $\beta_{5}$
$\beta_{6}$
$\alpha_{6}$
$\alpha_{5}$
2 copies
Figure 7. Two multicurves $\alpha$ and $\beta$ on the surface of genus three. There are to new components of $\alpha$ when compared to Figure 4: a nonseparating component (red) that we call $\alpha_{5}$ and a separating component (orange) that we call $\alpha_{6}$. Similarly, there are two new components of $\beta$: a separating component (blue) that we call $\beta_{5}$ and a nonseparating component (violet) that we call $\beta_{6}$.
We add the curves $\alpha_{5},\alpha_{6}$, two parallel copies of $\beta_{5}$ and finally the curve $\beta_{6}$. We note that $\beta_{5}$ and $\alpha_{6}$ are separating, and $\beta_{6}$ and $\alpha_{5}$ form a bounding pair. To read off the matrix $XX^{\top}$ for this pair of multicurves, we note two things. Firstly, the component $\alpha_{5}$ intersects the curves $\beta_{j}$ the same number of times as $\alpha_{1}$, except for the component $\beta_{5}$ of which there are two parallel copies. Secondly, the component $\alpha_{6}$ intersects each curve $\beta_{j}$ twice the number of times as $\alpha_{5}$, except for $\beta_{6}$. Using this, it is straightforward to calculate (ordering the curves as $\alpha_{6},\alpha_{5},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$)
$XX^{\top}=\begin{pmatrix}500&248&232&112&80&32\\ 248&124&116&56&40&16\\ 232&116&116&56&40&16\\ 112&56&56&28&20&8\\ 80&40&40&20&20&8\\ 32&16&16&8&8&4\end{pmatrix}$,
which is then checked to have irreducible characteristic polynomial.
5.2. Multicurve intersection degrees d<3g3d<3g-3
We now show how to modify our construction from Section 5.1 in order to realise multicurve intersection degrees smaller than the maximal multicurve intersection degree 3g33g-3. We need new building blocks to construct our surfaces.
Block 1. Our first block is obtained from the surface depicted in Figure 4, simply by dropping the component $\alpha_{3}$. A direct verification yields that for $y=1,2$ the characteristic polynomial of $XX^{\top}$ is irreducible and of degree 3.
Block 2. Our second block is obtained from the surface depicted in Figure 8. The characteristic polynomial of the matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$ is irreducible and of degree 1. Versions of this block with distinct characteristic polynomial can be obtained by taking $y$ parallel copies of $\beta$.
Figure 8. Two separating and filling curves $\alpha$ (red) and $\beta$ (blue) on the surface of genus two with one boundary component.
Block 3. Take a surface as depicted in Figure 9. We denote the red multicurve by $\alpha$ and the blue multicurve by $\beta$. The multicurve $\alpha$ has $k+1$ separating components: one for each of the handles that separates the handle, and one that separates all the handles. We denote the component of $\alpha$ that separates all the handles of the surface in Figure 9 by $\alpha_{1}$, and we denote the other separating components of $\alpha$ by $\alpha_{2},\alpha_{4},\dots,\alpha_{2k}$ from left to right. Finally, the remaining nonseparating components of $\alpha$ are $\alpha_{3},\alpha_{5},\dots,\alpha_{2k+1}$ from left to right.
$\cdots$
$\alpha_{1}$
$\alpha_{2}$
$\alpha_{2k}$
$\alpha_{3}$
$\alpha_{2k+1}$
$\beta_{1}$
$\beta_{k}$
$\beta_{0}$
Figure 9. A surface of genus $k$ with two boundary components, as well as two multicurves $\alpha$ (in red) and $\beta$ (in blue). The separating components of $\beta$ can have several parallel copies: the separating components $\beta_{1},\dots,\beta_{k}$ running through the handles have $y_{1},\dots,y_{k}$ copies, and the separating component $\beta_{0}$ to the right of the left-most handle has $y^{2}-k-4(y_{1}+\cdots+y_{k})$ copies.
The multicurve $\beta$ consists of a separating and a nonseparating component in each handle, as well as one separating curve $\beta_{0}$ to the right of the left-most handle. All the separating components of $\beta$ can have multiple parallel copies, see Figure 9. In this situation, we have
$XX^{\top}=\left(\begin{array}[]{c|ccccccc}4y^{2}&v_{y_{1}}^{\top}&v_{y_{2}}^{\top}&\cdots&v_{y_{k}}^{\top}\\ \hline\cr v_{y_{1}}&C_{y_{1}}&0&&\\ v_{y_{2}}&0&C_{y_{2}}&&\\ \vdots&&&\ddots&\\ v_{y_{k}}&&&&C_{y_{k}}\end{array}\right),~C_{y_{i}}=\begin{pmatrix}16y_{i}+4&8y_{i}\\ 8y_{i}&4y_{i}\end{pmatrix},~v_{y_{i}}=\begin{pmatrix}16y_{i}+4\\ 8y_{i}\end{pmatrix}$.
If all the $y_{i}$ are chosen pairwise distinct,  Remark 12 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. By Hilbert’s irreducibility theorem, there are infinitely many specifications of $y$ such that $y^{2}-k-4(y_{1}+\cdots+y_{k})>0$ and such that $\chi_{XX^{\top}}(t)\in\mathbb{Z}[t]$ is irreducible and of degree $2k+1$. Note that we can drop the separating components $\alpha_{2},\alpha_{4},\dots,\alpha_{2k}$ of $\alpha$ winding around one handle one by one in order to decrease the degree, reducing a 2-by-2 block to a 1-by-1 block, consisting of the coefficient $4y_{i}$, for each component dropped in this way. Again, if all the $y_{i}$ are chosen pairwise distinct, Remark 12 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. We can in this way construct all degrees $k+1\leq d\leq 2k+1$ for the surface of genus $k$ and 2 boundary components.
5.2.1. Realising multicurve intersection degrees $3g-6\leq d<3g-3$.
Using a block of type 1 or 2 instead of our standard starting surface depicted in Figure 4, we can reduce the multicurve intersection degree by 1 or 3, respectively. Since we use such a block on both sides of the surface in our construction, this gives the possibility to reduce the degree by any among the numbers 1,2,3,4 or 6. In particular, we can clearly realise the multicurve intersection degrees 3g4,3g53g-4,3g-5 and 3g63g-6. This argument works for $g\geq 4$.
In case of $g=3$, we need a separate argument. The idea is to copy our example of maximal degree from Figure 7, but leave out first $\alpha_{1}$ and then also $\alpha_{3}$. If we drop $\alpha_{1}$, which amounts do deleting the third row and column, we obtain a new matrix $XX^{\top}$ which is directly verified to have an irreducible characteristic polynomial. If we now drop also $\alpha_{3}$, which amounts to deleting the second-last row and column, we obtain a new matrix $XX^{\top}$ which again is verified to have irreducible characteristic polynomial. Since we have only dropped separating components, we have not changed the fact that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$.
We have therefore realised $d=5,4$ for $g=3$. For $g=3$, the case $d=3$ equals the case d=3g6d=3g-6, which is treated below.
5.2.2. Realising multicurve intersection degrees $g\leq d\leq 3g-6$.
We start by constructing a surface of genus g2g-2 with two boundary components, which we then close off in a second step.
Using surfaces of the type depicted in Figure 5 and applying the inductive step procedure illustrated in Figure 2, we can construct a surface of genus $g-2\geq 1$ and two boundary components, as well as filling multicurves $\alpha$ and $\beta$ with intersection degree 3(g2)1=3g73(g-2)-1=3g-7. The count of this intersection degree is exactly as in the non-Torelli case of Section 4, where we obtained degree 3g13g-1 on the closed surface of genus $g$ with 2 boundary components. Using at some point in the inductive procedure a block of type 3 of genus $k\leq g-2$, as depicted in Figure 9, we can reduce the degree by up to $2k-2\leq 2g-6$, realising multicurve intersection degrees from g1g-1 to 3g73g-7 on the surface of genus g2g-2 with two boundary components. Now we close the surface, as depicted in Figure 10, adding the new components $\alpha_{0}$ and $\beta_{0}$ to the multicurves $\alpha$ and $\beta$, respectively.
$\cdots$
$\beta_{0}$
$\alpha_{0}$
Figure 10. Two separating curves $\alpha_{0}$ and $\beta_{0}$. There are $\rho$ parallel copies of $\beta_{0}$.
We obtain the matrix
$XX^{\top}=\left(\begin{array}[]{c|ccc}64\rho+16a_{1}&4a_{1}&\dots&4a_{n}\\ \hline\cr 4a_{1}&&&\\ \vdots&&A&\\ 4a_{n}&&&\end{array}\right),$
where $A$ is the matrix $XX^{\top}$ before adding the curves $\alpha_{0}$ and $\beta_{0}$. Since $a_{1}=4a$, we can set $\rho=y^{2}-a$ to have the top-left coefficient $64y^{2}$, which is exactly the form of the matrix in 10. Finishing the argument as usual, we can realise the multicurve intersection degrees $g\leq d\leq 3g-6$ for $g\geq 3$.
5.2.3. Realising multicurve intersection degrees $1\leq d<g$.
Realising multicurve intersection degree one is clearly achieved by taking a pair of separating filling curves on the surface $S_{g}$.
For $2\leq d<g$, let us define f=g1df=g-1-d. We start with a surface block of type 3 of genus g2g-2, where we deleted all the components $\alpha_{2},\alpha_{4},\dots,\alpha_{2k}$ of $\alpha$. We also remove the component $\beta_{0}$ of $\beta$, see Figure 9. Furthermore, we let the $f+1\leq g-2$ first of the parameters $y_{i}$ be equal to $1$. Then we close off the surface as in the previous case, adding one component $\alpha_{0}$ to $\alpha$ and one component $\beta_{0}$ to $\beta$, compare with Figure 10. Assume there are $\rho$ parallel copies of $\beta_{0}$. We get
$XX^{\top}=\begin{pmatrix}64(\rho-g+2)+256\delta&32y_{1}&32y_{2}&\cdots&32y_{g-2}\\ 32y_{1}&4y_{1}&&&\\ 32y_{2}&&4y_{2}&&\\ \vdots&&&\ddots&\\ 32y_{g-2}&&&&4y_{g-2}\end{pmatrix}$,
where $\delta=y_{1}+\cdots+y_{g-2}$. We choose $\rho$ such that $64(\rho-g+2)+256\delta=64y^{2}$. To simplify the calculations, we let $z_{i}=4y_{i}$ for $i=1,\dots,g-2$. The matrix becomes
$XX^{\top}=\begin{pmatrix}64y^{2}&8z_{1}&8z_{2}&\cdots&8z_{g-2}\\ 8z_{1}&z_{1}&&&\\ 8z_{2}&&z_{2}&&\\ \vdots&&&\ddots&\\ 8z_{g-2}&&&&z_{g-2}\end{pmatrix}$.
By Lemma 9 in [LL24], the characteristic polynomial of $XX^{\top}$ equals
$p(t,y,\mathbf{z})=-64y^{2}\prod_{i=1}^{g-2}(t-z_{i})+t\prod_{i=1}^{g-2}(t-z_{i})-\sum_{i=1}^{g-2}64z_{i}^{2}\prod_{j\neq i}(t-z_{j}).$
If all the $z_{i}$ are pairwise distinct, this polynomial is irreducible as a polynomial in $t,y$ by Lemma 10 in [LL24]. However, we chose that the first $f+1$ coefficients $y_{1},\dots,y_{f+1}$ are equal to 11 and the other $y_{i}\neq 1$ and pairwise distinct. In particular, the polynomial $p(t,y)$ factors as $(t-4)^{f}\tilde{p}(t,y)$, where $\tilde{p}(t,y)$ is of degree g1f=dg-1-f=d in the variable $t$ and with pairwise distinct $z_{i}$. In particular, Lemma 10 in [LL24] implies that $\tilde{p}(t,y)\in\mathbb{Z}[t,y]$ is irreducible. Hilbert’s irreducibility theorem guarantees the existence of infinitely many positive specifications of $y$ such that the resulting polynomial is irreducible in $\mathbb{Z}[t]$.
This finishes the proof of 3.
Finally, we end this section with a proof of 5.
Proof of 5.
For every $g\geq 3$ and every integer $1\leq d\leq 3g-3$, we have constructed a filling pair of multicurves $\alpha$ and $\beta$, with a parameter $y$, such that $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t,y]$ is irreducible. By Remark 8, we may run the same argument to show that also the polynomial $\chi_{XX^{\top}}(t^{2},y)\in\mathbb{Z}[t,y]$ is irreducible. By Hilbert’s irreducibility theorem, we find infinitely many specifications of $y$ such that $\chi_{XX^{\top}}(t^{2})\in\mathbb{Z}[t]$ is irreducible of degree $2d$. The leading eigenvalue $\mu$ of $XX^{\top}$ is a root of a characteristic polynomial $\chi_{XX^{\top}}(t)$, so $\chi_{XX^{\top}}(t^{2})\in\mathbb{Z}[t]$ is the minimal polynomial of $\sqrt{\mu}$. Hence, the multicurve bipartite degree of $\alpha$ and $\beta$ equals $2d$.
For $g=2$, we use the example constructed in Section 4 for $d=3$ and the examples in [LL24] for $d=1,2$. Similarly to Remark 8, one can run the same proof as [LL24, Lemma 10] to show that $\chi_{XX^{\top}}(t^{2},y)\in\mathbb{Z}[t,y]$ is irreducible. ∎
6. Explicit pseudo-Anosov maps
There are two goals of this section, both inspired by a problem posed by Margalit [Mar19, Problem 10.4]. This section is completely independent of the proof of our main results, 1, 2 and 3.
The first goal is to construct, as explicitly as possible, pseudo-Anosov maps whose stretch factors have prescribed algebraic degrees. For suitably chosen examples, we can replace our nonsplitting criterion, 4, by the nonsplitting criterion of [LL24, Theorem 6]. This provides that the mapping class $T_{\alpha}\circ T_{\beta}$ has degree two over the trace field, which is more explicit than what we obtain using 4. However, we are still not able to control the exact number of parallel copies used for some of the components $\beta_{i}$. This construction is carried out in Section 6.1 and Section 6.2.
The second goal is to construct completely explicit pseudo-Anosov mapping classes of the form $T_{\alpha}\circ T_{\beta}$ with maximal stretch factor degree 6g66g-6. We are able to achieve this for $g\leq 201$ in Section 6.3.
6.1. Multicurve intersection degrees $d\leq 3g-3$
Recall that in Section 4, we constructed pairs of multicurves for which the Thurston–Veech construction realises the multicurve intersection degree 3g33g-3, for $g\geq 2$. In a first step, we build upon our examples from Section 4 to realise also all possible smaller multicurve intersection degrees $1\leq d<3g-3$ on $S_{g}$ for $g\geq 2$.
We need a new building block for our surfaces, which is depicted in Figure 11.
$y_{1}$
$y_{k}$
$\cdots$
$y^{2}-k$
$\alpha_{1}$
$\alpha_{2}$
$\alpha_{2k}$
$\alpha_{3}$
$\alpha_{2k+1}$
Figure 11. A surface of genus $k$ with two boundary components, as well as two multicurves $\alpha$ (in red) and $\beta$ (in blue). Some components of $\beta$ have several parallel copies, as indicated by $y_{1},\dots,y_{k}$ and $y^{2}-k$.
We denote the red multicurve by $\alpha$ and the blue multicurve by $\beta$. The multicurve $\alpha$ has $k+1$ separating components: one for each of the handles that separates the handle, and one that separates all the handles. We denote the component of $\alpha$ that separates all the handles of the surface in Figure 11 by $\alpha_{1}$, and we denote the other separating components of $\alpha$ by $\alpha_{2},\alpha_{4},\dots,\alpha_{2k}$ from left to right. Finally, the remaining nonseparating components of $\alpha$ are $\alpha_{3},\alpha_{5},\dots,\alpha_{2k+1}$ from left to right.
In this situation, we have
$XX^{\top}=\left(\begin{array}[]{c|ccccccc}4y^{2}&v^{\top}&v^{\top}&\cdots&v^{\top}\\ \hline\cr v&B_{y_{1}}&0&&\\ v&0&B_{y_{2}}&&\\ \vdots&&&\ddots&\\ v&&&&B_{y_{k}}\end{array}\right),\qquad B_{y_{i}}=\begin{pmatrix}4&2\\ 2&y_{i}\end{pmatrix},\qquad v=\begin{pmatrix}4\\ 2\end{pmatrix}$.
Let $p_{y_{i}}(t)=t^{2}-(4+y_{i})t+4(y_{i}-1)$ be the characteristic polynomial of $B_{y_{i}}$. We know from Section 4 that $p_{y_{i}}$ is irreducible if $y\geq 12$. So, choosing all $y_{i}\geq 12$ pairwise distinct, Remark 12 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. By Hilbert’s irreducibility theorem, there are infinitely many specifications of $y$ such that $y^{2}-k>0$ and such that $\chi_{XX^{\top}}(t)\in\mathbb{Z}[t]$ is irreducible and of degree $2k+1$.
Case 1: $2g\leq d<3g-3$. Assume we want to realise the multicurve intersection degree 3g3f3g-3-f for $0<f\leq g-3$. Let $k=f+2\leq g-1$. Start the inductive procedure as in Section 4 with the surface from Figure 11 as a starting point, adding g1kg-1-k more handles. The exact same argument yields a surface of genus g1g-1 with two boundary components, and a characteristic polynomial $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ that is irreducible and of degree 2k+1+3(g1k)=3g3k+12k+1+3(g-1-k)=3g-3-k+1. Closing up the surface exactly as in Section 4 yields 3g3k+2=3g3f3g-3-k+2=3g-3-f as a multicurve intersection degree on the closed orientable surface of genus $g$.
Case 2: $g<d<2g$. Assume we want to realise the multicurve intersection degree 2gf2g-f for $0<f\leq g-1$. Take the surface depicted in Figure 11 for k=g1k=g-1. Now, remove $f$ of the separating curve $\alpha_{2},\dots,\alpha_{2g-2}$. This slightly modifies the matrix $XX^{\top}$$f$ of the 2-by-2 blocks on the diagonal are now 1-by-1 blocks, with the single coefficient $y_{i}$. Nevertheless, since all the $y_{i}$ are chosen pairwise distinct, Remark 12 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. We note that for the coefficients $y_{i}$ in the 1-by-1 blocks, any positive integer can be chosen. By Hilbert’s irreducibility theorem, there are infinitely many specifications of $y$ such that $\chi_{XX^{\top}}(t)\in\mathbb{Z}[t]$ is irreducible and of degree 2g1f2g-1-f. Closing up the surface as in Section 4 yields the multicurve intersection degree 2gf2g-f on the closed orientable surface of genus $g$.
Case 3: $1\leq d\leq g$. This is the case we have already dealt with in [LL24].
6.2. Even degree stretch factors
In this section, we show that in our construction of multicurves in Section 6.1, the degree of the stretch factor of $T_{\alpha}\circ T_{\beta}$ equals two over the trace field. It uses the nonsplitting criterion of [LL24, Theorem 6] that we recall below.
Theorem 13 ([LL24], Theorem 6).
Let $\alpha,\beta\subset S$ be a pair of filling multicurves. Let $X$ be their geometric intersection matrix, let $d$ be their multicurve intersection degree and let $\Omega=\left(\begin{smallmatrix}0&X\\ X^{\top}&0\end{smallmatrix}\right)$. If $\dim(\Omega)>\sigma(\Omega+2I)+\mathrm{null}(\Omega+2I)>\dim(\Omega)-2d$, then the mapping class $T_{\alpha}\circ T_{\beta}$ is pseudo-Anosov with stretch factor $\lambda$ of degree $2d$.
Here, $\sigma(A)$ and $\mathrm{null}(A)$ denote the signature and the nullity, respectively, of the matrix $A$.
Theorem 14.
Let $\alpha$ and $\beta$ be an example of a pair of multicurves described in Section 6.1, realising a multicurve intersection degree $1\leq d\leq 3g-3$. Then the mapping class $T_{\alpha}\circ T_{\beta}$ is pseudo-Anosov with stretch factor $\lambda$ of degree $2d$.
For the case $1\leq d\leq g$, this is shown in [LL24].
Proof of 14.
According to 13, all there is to show is
(1) $\dim(\Omega)>\sigma(\Omega+2I)+\mathrm{null}(\Omega+2I)>\dim(\Omega)-2d.$
We now make a case distinction depending on $d$.
Case 1: $2g\leq d\leq 3g-3$. We consider the submatrix $\Omega^{\prime}$ of $\Omega$ that is obtained by deleting all the rows and columns corresponding to components of the multicurve $\alpha$ that have been added during the inductive step or closing up of the surface. Furthermore, if d<3g3d<3g-3, we also remove the component of $\alpha$ encircling multiple handles of the starting surface, that is, the component $\alpha_{1}$ of the surface depicted in Figure 11.
A base change by a permutation matrix brings $\Omega^{\prime}+2I$ into block diagonal form with g1g-1 blocks corresponding to genus one surface pieces as depicted in Figure 1, and a block of the form $2I$. For a block of the former type, and for $y>4$, we directly calculate that the nullity is zero and the signature equals the dimension of the block minus two. Already, this implies that certainly the signature of $\Omega+2I$ is not equal to its dimension, and it only remains to verify the lower bound in Equation 1.
By construction, if the genus equals $g\geq 2$, we have g1g-1 surface pieces as in Figure 1. This in particular implies that $\sigma(\Omega^{\prime})=\dim(\Omega^{\prime})-2g+2$. Furthermore, we have $\dim(\Omega)-\dim(\Omega^{\prime})=d-2g+2$. The latter equality follows from that fact that the number of components of $\alpha$ in our construction is exactly $d$, and there are two components per surface pieces as in Figure 1. We now calculate
$\sigma(\Omega+2I)$ $\geq\sigma(\Omega^{\prime})-(\dim(\Omega)-\dim(\Omega^{\prime}))$
$=(\dim(\Omega^{\prime})-2g+2)-(d-2g+2)$
$=\dim(\Omega^{\prime})-d$
$>\dim(\Omega)-2d,$
which implies Equation 1, so we are done for this case.
Case 2: $g<d<2g$. We consider the submatrix $\Omega^{\prime}$ of $\Omega$ that is obtained by deleting two rows and two columns corresponding to the following components of the multicurve $\alpha$: the one corresponding to the component $\alpha_{1}$ encircling multiple handles in Figure 11 and the one obtained from closing the surface. Recall that we have removed f=2gdf=2g-d of the separating curves $\alpha_{2},\dots,\alpha_{2g-2}$.
A base change by a permutation matrix brings $\Omega^{\prime}+2I$ into block diagonal form with g1fg-1-f blocks corresponding to surface pieces as in Figure 1ff blocks corresponding to surface pieces as in Figure 1 but with the separating component of $\alpha$ removed, and a block of the form $2I$.
For a block of the first type, and for $y>4$, recall from the previous case that the nullity is zero and the signature equals the dimension of the block minus two. For a block of the second type, the sum of the nullity and the signature equals the dimension of the block if $y\leq 3$, and it equals the dimension of the block minus two if $y>3$. We may assume that for at least one block of the second type, we have $y=3$. This is enough to ensure that $\dim(\Omega)>\sigma(\Omega+2I)+\mathrm{null}(\Omega+2I)$, so again we only need to verify the lower bound in Equation 1.
By construction, if the genus equals $g\geq 2$, we have g1g-1 surface pieces as in Figure 1. Having at least one piece with $y\leq 3$, this implies that $\sigma(\Omega^{\prime})>\dim(\Omega^{\prime})-2g+2$. Furthermore, we have $\dim(\Omega)-\dim(\Omega^{\prime})=2$. We now calculate
$\sigma(\Omega+2I)$ $\geq\sigma(\Omega^{\prime})-(\dim(\Omega)-\dim(\Omega^{\prime}))$
$>(\dim(\Omega^{\prime})-2g+2)-2$
$=\dim(\Omega^{\prime})-2g$
$=\dim(\Omega)-2g+2\geq\dim(\Omega)-2d,$
which implies Equation 1 also in the case $g<d<2g$, so we are done. ∎
6.3. Explicit examples with stretch factor degree 6g66g-6
We conclude this section by giving explicit computations, for $1<g\leq 201$, supporting a conjecture on the irreducibility of the characteristic polynomials constructed in Section 4 for specific values of $y$. In the inductive step of Section 4, one uses a map
$\phi_{k}:\begin{array}[]{cccc}M_{k}(\mathbb{Z})\times\mathbb{Z}&\longrightarrow&M_{k+3}(\mathbb{Z})\\ (C,y)&\mapsto&\left(\begin{array}[]{c|c|c}4y^{2}&*&*\\ \hline\cr*&C&\\ \hline\cr*&&A\end{array}\right)\end{array},\qquad\mathrm{with}\ A=\left(\begin{smallmatrix}4&2\\ 2&12\end{smallmatrix}\right).$
For $g>1$ we inductively construct the $(3g-1)\times(3g-1)$ matrix $M_{g}$ with the maps $\phi_{3i-1}$ for $i=1,\dots,g-1$:
$M_{g}=\phi_{3(g-1)-1}(\phi_{3(g-2)-1}(\dots\phi_{3\cdot 2-1}(\phi_{3\cdot 1-1}(B,y^{(1)}),y^{(2)}),\dots,y^{(g-2)}),y^{(g-1)}),$
with $B=\left(\begin{smallmatrix}4&2\\ 2&13\end{smallmatrix}\right)$ and suitable parameters $y^{(i)}$ given by Hilbert’s irreducibility theorem. The condition $y^{2}>\frac{1}{4}c_{11}+1$ appearing in the construction is obviously equivalent to $(y^{(i)})^{2}>(y^{(i-1)})^{2}+1$. Finally, following Section 4 the matrix $XX^{\top}$ for the multicurves $\alpha$ and $\beta$ on the closed surface of genus $g+1$ takes the form
$N_{g}=\left(\begin{array}[]{c|ccc}y^{2}&*\\ \hline\cr*&M_{g}\end{array}\right)$
with the condition $y^{2}>\frac{1}{4}(M_{g})_{11}=(y^{(g-1)})^{2}$.
By computer assistance [S+YY], one immediately checks the following proposition.
Proposition 15.
For any $1<g\leq 200$, if $y^{(i)}=i+1$ for $i=1,\dots,g-1$, then the characteristic polynomial $\chi_{M_{g}}$ is irreducible over $\mathbb{Q}$. Moreover for $y=g+1$, $\chi_{N_{g}}$ is irreducible over $\mathbb{Q}$.
Together with 14, this gives explicit examples of pseudo-Anosov maps realising the upper bound 6g66g-6 in 3 for every $1<g\leq 201$. We do not know whether $\chi_{M_{g}}$ and $\chi_{N_{g}}$ are actually irreducible for every $g>200$ with the parameters $y^{(i)}=i+1$ chosen as in 15.